A skilled jet fighter flies a stunt plane in a vertical circle of 1200 ft
radius at constant speed of 140 mi/h1. Calculate the jet fighter’s
acceleration at
(i) the highest point
(ii) the lowest point
1
Expert's answer
2021-09-15T11:38:13-0400

(i) At the highest point:


"a=g-\frac{v^2}r=32-\frac{205^2}{1200}=-3\text{ ft/s}^2\text{ (acts upward)}."

(ii) At the lowest point


"a=g-\frac{v^2}r=32+\frac{205^2}{1200}=67\text{ ft/s}^2\text{ (downward)}."